OpenStudy (anonymous):

The functions f and g are differentiable and f(g(x))=x^2 for all x. If f(4)=8, g(4)=8, f'(8)=-2, what is the value of g'(4)?

jimthompson5910 (jim_thompson5910):

f(g(x))=x^2 f'(g(x))*g'(x)=2x f'(g(4))*g'(4)=2(4) Do you see what to do from here?

OpenStudy (anonymous):

What should I do next?

jimthompson5910 (jim_thompson5910):

g(4) = ??

OpenStudy (anonymous):

I don't get it.

jimthompson5910 (jim_thompson5910):

what is g(4)

OpenStudy (anonymous):

8.

jimthompson5910 (jim_thompson5910):

so f'(g(4))*g'(4)=2(4) becomes f'(8)*g'(4)=2(4)

jimthompson5910 (jim_thompson5910):

what is f'(8)

OpenStudy (anonymous):

-2.

jimthompson5910 (jim_thompson5910):

now replace f'(8) with -2 and solve for g'(4)

OpenStudy (anonymous):

So the answer is -2?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

Thanks for the help. Sorry I was clueless.